OpenStudy (anonymous):

The score on the statistics class final exam follow normal distribution with a mean of 80 and a standard deviation of 5. What percent of students scored 90 or better ?

OpenStudy (anonymous):

\[P(X\ge90)=P\left(\frac{X-80}{5}\ge\frac{90-80}{5}\right)=P(Z\ge2)\] You can use this rule to quickly compute the proportion: http://en.wikipedia.org/wiki/68%E2%80%9395%E2%80%9399.7_rule